OpenStudy (anonymous):

Help please? Solve the absolute value equation. Graph the solution. 2|3x – 5| – 8 = 8

OpenStudy (anonymous):

2|3x – 5| – 8 = 8 or 2|3x – 5| – 8 = -8

OpenStudy (anonymous):

how would i turn these into fractions? @zordoloom

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

2|3x – 5| – 8 = 8 add 8 2|3x – 5|=16 divide by 2 |3x – 5|=8 now, 3x-5=-8 or 3x-5=8 solve the first one: 3x-5=-8 add 5 3x=-3 x=-1 now solve the second one: 3x-5=8 add 5 3x=13 now divide by 3 x=13/3 so, x=13/3 or x=-1

OpenStudy (anonymous):

sorry! i was eating lunch...

OpenStudy (anonymous):

thank you!! i read through it all ^^